A company has an annual coupon bond issue that has a coupon rate of 7%, a par value of $1,000, and a current price of $1,153.19. Determine the bond’s YTC if the bond is called back 4 years from now with a call premium of 10%.

Group of answer choices

5% 10% 7% $1,100

A $1,000 par value bond has an 8% coupon rate (paid semiannually). It has 5 years remaining to maturity. If bond’s current price $1,085.30, what should be the YTM of this bond?

7% 6% 5% 4%

A $1,000 par value bond has an 8% coupon rate (paid semiannually) and a current price of $1,148.77. What should be the YTC of this bond if the bond is called back 5 years from now with a call price of $1047.20?  

6.8% 3.5% 5.4% 2.9%

What is the price of a semiannual coupon bond 1 year from now if the bond has a current maturity of 10 years, a coupon rate of 8%, and a par value $1,000? The market interest rate is 6%.

$1,085.30

$1,000

$1,231.15

$1,137.54
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

SOLVED WITH BA II PLUS CALCULATOR

C w d ENG 00:17 30-07-2020 99+ 31 Х ET196 ER ES ET EU EV EW EX EY EZ FA FB FC FD 174 175 176 177 178 179 180 BA II PLUS first

C w d ENG 00:19 30-07-2020 99+ 31 Х S140 . L M N O Р Q R S T U V BA II PLUS calculator strokes first of all set 2 C/Y = P/Y =

C w d ENG 00:21 30-07-2020 99+ 31 Х LD 20 КХ KY KZ LA LB LC LD LE LE 1 2 3 4 5 6 calculator strokes FOR YTC 7 8 first of all

C w d ENG 00:22 30-07-2020 99+ 31 Х 0222 fix . L M N Р Q R S T U V 205 206 BA II PLUS price of semi annual coupon bond set 2